Which of the following is part of the set of all possible rational zeros of



f(x)=2x4−x3+6x2−5x−8?



1
Expert's answer
2022-06-08T17:51:48-0400

Since all coefficients are integers, we can apply the rational zeros theorem.

The trailing coefficient (the coefficient of the constant term) is "-8."

Find its factors: "\pm1, \pm2, \pm4,\pm8."

These are the possible values for "p."

The leading coefficient is "2."

Find its factors: "\pm1, \pm2."

These are the possible values for "q."

Find all possible values of "f/q: \pm\dfrac{1}{1}, \pm\dfrac{1}{2},\pm\dfrac{2}{1},\pm\dfrac{2}{2},\pm\dfrac{4}{1},\pm\dfrac{4}{2},\pm\dfrac{8}{1},\pm\dfrac{8}{2}."

Simplify and remove the duplicates (if any).

These are the possible rational roots: "\pm1, \pm\dfrac{1}{2}, \pm2, \pm4,\pm8."